# Three-dimensional Abelian and nonAbelian gauge Higgs theories

### v1

@@ 2.1 Phase transitions and critical phenomena
Classical and quantum phase transitions are phenomena of great interest in modern physics, both theoretically and experimentally. Classical phase transitions are generally driven by thermal fluctuations [5, 112, 94, 113, 114, 6, 115, 2, 95, 116, 117], while their quantum counterparts arise from quantum fluctuations in the zero-temperature limit [118, 14, 101, 117]. Phase transitions separate different phases characterized by distinctive properties. In many cases they are associated with the spontaneous breaking of a global symmetry, due to the condensation of an order parameter and the emergence of a corresponding ordering. Notable exceptions occur in the presence of gauge symmetries. In this case the transition may be driven by topological excitations [119, 12, 120, 20], without a local order parameter and breaking of a global symmetry.
@@
Paradigmatic examples of models undergoing such finite-temperature continuous transitions are the $N$-vector models, defined by the Hamiltonian
@@
In this review we discuss statistical systems undergoing classical transitions. However, many results on the general features of the transitions and the corresponding critical behavior can be extended to quantum transitions in a lower dimension, according to the above-mentioned quantum-to-classical mapping.
@@ 2.2 The field-theoretical Landau-Ginzburg-Wilson approach to continuous phase transitions
    The scaling hypothesis, which assumes that the singular part of the thermodynamic properties and of the correlation functions of the order parameter are scale invariant at criticality. Close to the critical point, they satisfy scaling relations when expressed in terms of appropriately defined scaling fields.
@@
We now present more details of some specific examples of the LGW approach that will be particularly relevant in the following of this review.
@@ 2.2.1 O($N$)-symmetric $\Phi^{4}$ theory
The simplest example is the O($N$)-symmetric $\Phi^{4}$ theory, defined by the Lagrangian density
@@ 2.2.2 U($N$)-symmetric $\Phi^{4}$ theory
We now discuss a straightforward generalization of the O($N$) $\Phi^{4}$ theory, appropriate to describe transitions in U($N$)-symmetric models, driven by a complex $N$-dimensional vector order parameter. The corresponding LGW theory is constructed using a complex $N$-component vector $\Psi^{a}$ as order-parameter field. Then, as one can easily check, the most general U($N$)-symmetric LGW Lagrangian requires only one quartic term, i.e.
@@
A remarkable property of this U($N$)-invariant LGW theory is that it is actually symmetric under a larger O($2N$) symmetry group. This can be easily seen by noting that the Lagrangian (3) can be exactly mapped into that of the O($2N$)-vector models, cf. Eq. (2), by rewriting $\Psi^{a}$ in terms of real components, i.e. $\Psi^{a}=\Phi^{a}+i\Phi^{a+N}$, where $\Phi^{a}$ is a real $(2N)$-component vector. As one can easily check, U($N$)-symmetric terms that break the O($2N$) symmetry require higher powers of the field ${\bm{\Psi}}$. Indeed the lowest-dimensional scalar operator breaking the O($2N$) symmetry is the dimension-six operator $\sum_{\mu}|\bar{\bm{\Psi}}\cdot\partial_{\mu}{\bm{\Psi}}|^{2}$.
Therefore, the LGW theory predicts that the critical behaviors of U($N$)-symmetric systems, at phase transitions driven by the condensation of a complex $N$-component order parameter, belong to the O($2N$) vector universality classes, which implies an effective enlargement of the symmetry of the critical modes, from U($N$) to O($2N$). The U($N$)-symmetric RG perturbations breaking the O($2N$) symmetry are irrelevant in the critical limit, giving only rise to suppressed scaling corrections.
@@ 2.2.3 Generic Landau-Ginzburg-Wilson $\Phi^{4}$ theory with a single quadratic invariant
There are physically interesting transitions that are described by more general LGW $\Phi^{4}$ field theories and are characterized by more complex symmetries and symmetry-breaking patterns, see, e.g., Refs. [147, 130, 2, 95, 148, 149]. The most general LGW Lagrangian for a $N$-component order-parameter field $\Phi_{i}({\bm{x}})$ can be written as
@@
The number of independent quadratic $r_{i}$ and quartic $u_{ijkl}$ parameters crucially depends on the symmetry group of the theory. An interesting class of models are those in which $\sum_{i}\Phi^{2}_{i}$ is the unique quadratic term that is invariant under the symmetry group of the theory [147, 2]. In this case, all $r_{i}$ are equal, $r_{i}=r$, and $u_{ijkl}$ must satisfy the trace condition [147] $\sum_{i}u_{iikl}\propto\delta_{kl}$, to avoid the generation of other quadratic invariant terms under RG transformations. In these models, criticality is driven by tuning the single parameter $r$, which may correspond to the reduced temperature. Therefore, these cases correspond to the prototypical transitions outlined in Sec. 2.1. All field components become critical simultaneously and the two-point function in the disordered phase is diagonal, that is, $G_{ij}({\bm{x}},{\bm{y}})\equiv\langle\Phi_{i}({\bm{x}})\Phi_{j}({\bm{y}})\rangle=\delta_{ij}G({\bm{x}},{\bm{y}})$. LGW theories with multiparameter quartic terms provide effective descriptions of complex phase transitions in several physical contexts [95]. We mention anisotropic magnets  [130, 150], disordered systems [151, 152], frustrated systems [153, 154, 155, 156, 157, 158], spin and density wave models [159, 160, 161, 162, 163, 164], LAH models in the strong gauge-coupling limit [165, 166, 84], multiflavor scalar chromodynamics [97, 167], and also the finite-temperature chiral transition in hadronic matter with $N_{f}$ light flavors [25, 168, 169].
The RG flow in $\Phi^{4}$ theories, such as those with Lagrangian density (4), has generally several FPs. Among them, the stable FP controls the critical behavior of the corresponding continuous transitions, and therefore can be associated with their universality class. If there is no stable FP, the transitions in the corresponding class of models are generally expected to be first order, unless the considered field-theoretical model misses some relevant features of the transitions under investigation.
@@
LGW theories (4) in which the symmetry allows several quadratic terms are also relevant phenomenologically. In particular, they provide effective field-theoretical models describing the multicritical behaviors that occur in the presence of competing orderings [170, 171, 172]. The main features of the critical behaviors at multicritical points (MCPs) are discussed in A.4.
As a paradigmatic example, one may consider multicritical LGW theories that describe the competition of order parameters associated with two different global O($N_{1}$) and O($N_{2}$) symmetries. Their Lagrangian density is given by [170, 171, 172, 95, 173]
${\cal L}_{\rm mc}=\sum_{a=1}^{2}\Bigl{[}\partial_{\mu}{\bm{\phi}}_{a}\cdot\partial_{\mu}{\bm{\phi}}_{a}+r_{a}\,{\bm{\phi}}_{a}\cdot{\bm{\phi}}_{a}+v_{a}\,({\bm{\phi}}_{a}\cdot{\bm{\phi}}_{a})^{2}\Bigr{]}+w\,{\bm{\phi}}_{1}\cdot{\bm{\phi}}_{1}\,{\bm{\phi}}_{2}\cdot{\bm{\phi}}_{2},$ (5)
@@
The phase diagram of the model (5) shows different disordered and ordered phases, separated by transition lines meeting at one point. A standard mean-field analysis of the model shows that, by varying the quartic coupligs $v_{1}$, $v_{2}$, and $w$, one can obtain qualitatively different phase diagrams [170, 171, 172], as shown in Fig. 1. The multicritical behavior of models associated with a multicritical LGW theory is again determined by the RG flow in the space of the quartic parameters [170, 171, 172, 95, 173, 175, 176, 174, 177, 178]. Again, a continuous multicritical behavior is generally associated with a stable stable FP of such RG flow.
@@ 3 Phase transitions and critical phenomena in the presence of gauge symmetries
The description of the main features of phase transitions in the presence of gauge symmetries is more complex than that outlined in the previous Section. First of all, while conventional transitions are always related with the condensation of a local order parameter, in the presence of gauge fields, we may have transitions without local order parameter (topological transitions). When an order parameter is present, we also need to distinguish its behavior under gauge transformations: the order parameter may be gauge invariant as well as not gauge invariant. Moreover, the role of gauge fields may differ. In some transitions, gauge fields are only relevant to define the set of critical modes (the gauge-invariant ones), while in some other cases, the phase transition is due to a nontrivial interplay between matter and gauge fields, so that gauge fields determine the critical behavior and should be included in the corresponding effective field theory (these transitions will be named charged transitions).
It is worth noting that the study of the critical behavior of gauge-symmetric models is also relevant to understand the critical behavior of unconventional transitions in condensed-matter physics, in which a gauge symmetry is supposed to emerge at criticality, see, e.g., Refs.  [7, 12, 13, 15, 179, 180, 20, 21]. Unconventional quantum transitions [181, 20, 21] are observed between topological phases [182, 183, 184, 69], in the context of the integer and fractional quantum Hall effect [185, 186, 187, 188]. We also mention transitions between topologically-ordered and spin-ordered phases, between topologically-ordered and valence-bond-solid (VBS) phases [27], and direct transitions between differently ordered phases, such as the Néel-to-VBS transition of 2D quantum magnets [33, 189]. Quantum gauge-field theories provide effective descriptions of deconfined quantum transitions that exhibit fractionalization of the low-energy excitations, topological order, and long-range entanglement. Indeed, they admit an effective description in terms of field theories with emergent gauge fields coupled to matter fields carrying fractional quantum numbers of the microscopic global symmetry, see, e.g., Refs. [20, 21] for reviews. Studies of deconfined transitions are reported in, e.g., Refs. [33, 189, 74, 72, 34, 37, 190, 191, 192, 39, 38, 40, 80, 31, 32, 83, 193, 194, 195, 196, 197, 198, 199, 200, 201, 21], and references therein.
@@ 3.1 Different types of critical scenarios in statistical models with gauge symmetries
As we have mentioned above, in the presence of gauge symmetries, the effective LGW approach discussed in Sec. 2 cannot be generically applied. Other approaches should also be used to obtain an effective description of the critical behavior. The choice of the approach to be used depends on the main features of the mechanism driving the particular phase transition of the lattice gauge model under study.
@@
    LGW transitions, in which only gauge-invariant correlations of the matter field are critical. They admit a conventional effective LGW field-theoretical description in terms of a gauge-invariant order-parameter field, which orders at the transition, driving the spontaneous breaking of the global symmetry. Gauge fields do not display critical correlations. The gauge symmetry only prevents non-gauge-invariant observables from becoming critical.
@@
    ${}^{\,\times}$ transitions, in which only matter fields develop critical correlations, similarly to LGW transitions, however the local order parameter driving the spontaneous symmetry breaking at the transition is not gauge invariant. They also admit an LGW description, although in this case the order-parameter field does not correspond to gauge-invariant modes. [^3] It is important to note that gauge symmetry breaking never occurs [202, 203, 43, 8], despite the common use of referring to the Higgs mechanism as spontaneous gauge symmetry breaking. The statement that in some cases a non-gauge invariant order parameter is needed does not refer to the description of the thermodynamic phases, but to the construction of the effective model which encodes the universal critical properties of a continuous phase transition. This means that, in some cases, the critical behavior of gauge-invariant quantities can only be correctly described by assuming that it is inherited from a more fundamental gauge-dependent order parameter, which can in principle be defined, e.g., by using an appropriate gauge fixing condition. Therefore, to distinguish them from standard LGW transitions, we added the $\times$ superscript.  [^4] Note that transitions with similar properties have been sometimes called LGW transitions, see, e.g., [21]. In particular, systems with discrete gauge groups undergo LGW× transitions for sufficiently weak gauge couplings.
@@
    GFT transitions, in which the critical behavior is due to a nontrivial interplay of critical modes arising from both matter and gauge fields. Therefore, their description requires an effective gauge field theory (GFT), with dynamical gauge and matter fields. GFT transitions have been observed in the presence of continuous Abelian and non-Abelian gauge symmetries. They also occur in models that are locally invariant under discrete groups when the gauge symmetry enlarges at criticality, so that the critical modes are invariant under a continuous group.
@@
    Topological transitions, which are driven by topological modes, in which one cannot identify a local order-parameter field. A typical example is provided by the topological transition of the 3D lattice ${\mathbb{Z}}_{2}$-gauge model [119].

### v3

@@ 2.1 Phase transitions and critical phenomena
Classical and quantum phase transitions are phenomena of great interest in modern physics, both theoretically and experimentally. Classical phase transitions are generally driven by thermal fluctuations [5, 113, 95, 114, 115, 6, 116, 2, 96, 117, 118], while their quantum counterparts arise from quantum fluctuations in the zero-temperature limit [119, 14, 102, 118]. Phase transitions separate different phases characterized by distinctive properties. In many cases they are associated with the spontaneous breaking of a global symmetry, due to the condensation of an order parameter and the emergence of a corresponding ordering. Notable exceptions occur in the presence of gauge symmetries. In this case the transition may be driven by topological excitations [120, 12, 121, 20], without a local order parameter and without breaking any global symmetry.
@@
Paradigmatic examples of models undergoing finite-temperature continuous transitions are the $N$-vector models, defined by the Hamiltonian
@@
In this review we discuss statistical systems undergoing classical transitions. However, many results on the general features of the transitions and the corresponding critical behavior can be extended to quantum transitions in a lower dimension by means of the quantum-to-classical mapping.
@@ 2.2 The field-theoretical Landau-Ginzburg-Wilson approach to continuous phase transitions
    The scaling hypothesis, which assumes that the singular part of the thermodynamic observables and of the correlation functions of the order parameter are scale invariant at criticality. Close to the critical point, they satisfy scaling relations when expressed in terms of appropriately defined scaling fields.
@@
We now present more details for some specific LGW theories that will be particularly relevant in the following.
@@ 2.2.1 O($N$)-symmetric $\Phi^{4}$ theory
The simplest LGW model is the O($N$)-symmetric $\Phi^{4}$ theory, defined by the Lagrangian density
@@ 2.2.2 U($N$)-symmetric $\Phi^{4}$ theory
We now discuss a straightforward generalization of the O($N$) $\Phi^{4}$ theory, appropriate to describe transitions in U($N$)-symmetric models driven by a complex $N$-dimensional vector order parameter. The corresponding LGW theory is defined in terms of a complex $N$-component vector field ${\bm{\Psi}}({\bm{x}})$. The most general U($N$)-symmetric LGW Lagrangian is
@@
A remarkable property of the U($N$)-invariant LGW Lagrangian is that it is symmetric under a larger O($2N$) symmetry group. This can be easily seen by noting that the Lagrangian (3) is equivalent to that of the O($2N$)-vector models, reported in Eq. (2), if the complex field component $\Psi^{a}$ is rewritten as $\Psi^{a}=\Phi^{a}+i\Phi^{a+N}$, where $\Phi^{a}$ is a real $(2N)$-component vector. As one can easily check, U($N$)-symmetric terms that break the O($2N$) symmetry require higher powers of the field ${\bm{\Psi}}$. Indeed, the lowest-dimensional scalar operator breaking the O($2N$) symmetry is the dimension-six operator $\sum_{\mu}|\bar{\bm{\Psi}}\cdot\partial_{\mu}{\bm{\Psi}}|^{2}$.
Therefore, the LGW theory predicts that the critical behavior of U($N$)-symmetric systems, at phase transitions driven by the condensation of a complex $N$-component order parameter, belong to the O($2N$) vector universality class, which implies an effective enlargement of the symmetry of the critical modes, from U($N$) to O($2N$). The U($N$)-symmetric RG perturbations breaking the O($2N$) symmetry are irrelevant in the critical limit, giving only rise to suppressed scaling corrections.
@@ 2.2.3 Generic Landau-Ginzburg-Wilson $\Phi^{4}$ theory with a single quadratic invariant
Beside O($N$) and U($N$) invariant models, there are physically interesting transitions that are described by LGW $\Phi^{4}$ field theories characterized by more complex symmetries and symmetry-breaking patterns, see, e.g., Refs. [176, 131, 2, 96, 177, 178]. The most general LGW Lagrangian for an $N$-component order-parameter field $\Phi_{i}({\bm{x}})$ (with $i=1,2,...N$) can be written as
@@
The number of independent quadratic $r_{i}$ and quartic $u_{ijkl}$ parameters crucially depends on the symmetry group of the theory. An interesting class of models are those in which $\sum_{i}\Phi^{2}_{i}$ is the unique quadratic term that is invariant under the symmetry group of the theory [176, 2]. In this case, all $r_{i}$ are equal, $r_{i}=r$, and $u_{ijkl}$ must satisfy the trace condition [176] $\sum_{i}u_{iikl}\propto\delta_{kl}$, to avoid the generation of other quadratic invariant terms under RG transformations. In these models, criticality is driven by tuning the single parameter $r$, which may correspond to the reduced temperature. Therefore, these cases correspond to the prototypical transitions outlined in Sec. 2.1. All field components become critical simultaneously and the two-point function in the disordered phase is diagonal, that is, $G_{ij}({\bm{x}},{\bm{y}})\equiv\langle\Phi_{i}({\bm{x}})\Phi_{j}({\bm{y}})\rangle=\delta_{ij}G({\bm{x}},{\bm{y}})$. LGW theories with multiparameter quartic terms provide effective descriptions of complex phase transitions in several physical contexts [96]. We mention anisotropic magnets (see, e.g., Refs. [131, 179]), disordered systems (see, e.g., Refs. [180, 181, 182, 183]), frustrated systems (see, e.g., Refs. [184, 185, 186, 187, 188, 189]), spin and density wave models (see, e.g., Refs. [190, 191, 192, 193, 194, 195]), LAH models in the strong gauge-coupling limit (see, e.g., Refs. [196, 197, 84]), multiflavor scalar chromodynamics [98, 198], and also the finite-temperature chiral transition in hadronic matter with $N_{f}$ light flavors [25, 199, 200].
The RG flow in $\Phi^{4}$ theories, such as those with Lagrangian density (4), has generally several FPs. Among them, the stable FP controls the critical behavior of the corresponding continuous transitions, and therefore can be associated with their universality class. If there is no stable FP, the transitions in the corresponding class of models are generally expected to be of first order, unless the considered field-theoretical model misses some relevant features of the transitions under investigation.
@@
LGW theories (4) in which the symmetry allows several quadratic terms are also relevant phenomenologically. In particular, they provide effective field-theoretical models describing the multicritical behaviors that occur in the presence of competing orderings [201, 202, 203]. The main features of the critical behavior at multicritical points (MCPs) are discussed in A.4.
@@
${\cal L}_{\rm mc}=\sum_{a=1}^{2}\Bigl{[}\partial_{\mu}{\bm{\phi}}_{a}\cdot\partial_{\mu}{\bm{\phi}}_{a}+r_{a}\,{\bm{\phi}}_{a}\cdot{\bm{\phi}}_{a}+v_{a}\,({\bm{\phi}}_{a}\cdot{\bm{\phi}}_{a})^{2}\Bigr{]}+w\,{\bm{\phi}}_{1}\cdot{\bm{\phi}}_{1}\,{\bm{\phi}}_{2}\cdot{\bm{\phi}}_{2}.$ (5)
[…]
The phase diagram of the model (5) shows different disordered and ordered phases, separated by transition lines meeting in one point. A standard mean-field analysis of the model shows that, by varying the quartic coupligs $v_{1}$, $v_{2}$, and $w$, one can obtain qualitatively different phase diagrams [201, 202, 203], as shown in Fig. 1. The multicritical behavior of models associated with a multicritical LGW theory is again determined by the RG flow in the space of the quartic parameters [201, 202, 203, 96, 204, 206, 152, 205, 207, 208]. A multicritical universality class is generally associated with a stable FP of such RG flow.
@@ 3 Phase transitions and critical phenomena in the presence of gauge symmetries
The description of the main features of phase transitions in the presence of gauge symmetries is more complex than that outlined in the previous Section. First of all, while conventional transitions are always related to the condensation of a local order parameter, in the presence of gauge fields, we may have transitions without local order parameter (topological transitions). When an order parameter is present, we also need to distinguish its behavior under gauge transformations: the order parameter may be gauge invariant as well as not gauge invariant. Moreover, the role of gauge fields may differ. In some transitions, gauge fields are only relevant to define the set of critical modes (the gauge-invariant ones), while in some other cases, the phase transition is due to a nontrivial interplay between matter and gauge fields, so that gauge fields determine the critical behavior and should be included in the corresponding effective field theory (these transitions will be named charged transitions).
The effective description of the different critical behaviors developed by systems with gauge symmetries requires different effective approaches, such as the standard gauge-invariant LGW approach when gauge fields are not critical, or the more complex gauge field-theory approach with explicit gauge fields in the presence of critical gauge correlations. This distinction is crucial to understand which statistical field theory is actually realized by the various continuous phase transitions occurring in the presence of gauge symmetries.
It is worth noting that the study of the critical behavior of gauge-symmetric models is also relevant to understand the critical behavior of unconventional transitions in condensed-matter physics, in which a gauge symmetry is supposed to emerge at criticality, see, e.g., Refs.  [7, 12, 13, 15, 209, 210, 20, 21]. Unconventional quantum transitions [211, 20, 21] are observed between topological phases [212, 213, 214, 69], in the context of the integer and fractional quantum Hall effect [215, 216, 217, 218]. We also mention transitions between topologically-ordered and spin-ordered phases, between topologically-ordered and valence-bond-solid (VBS) phases [27], and direct transitions between differently ordered phases, such as the Néel-to-VBS transition of 2D quantum magnets [33, 219]. Deconfined quantum transitions that exhibit fractionalization of the low-energy excitations, topological order, and long-range entanglement also admit an effective description in terms of field theories with emergent gauge fields coupled with matter fields carrying fractional quantum numbers of the microscopic global symmetry, see, e.g., Refs. [20, 21] for reviews. Studies of deconfined transitions are reported in, e.g., Refs. [33, 219, 74, 72, 34, 37, 220, 221, 222, 39, 38, 40, 80, 31, 32, 83, 223, 224, 225, 226, 227, 228, 229, 230, 21], and references therein.
@@ 3.1 Different types of critical scenarios in statistical models with gauge symmetries
As we have mentioned above, in the presence of gauge symmetries, the effective LGW approach discussed in Sec. 2 cannot be generically applied. Other approaches should also be used to obtain an effective description of the critical behavior. The choice of the appropriate approach depends on the main features of the mechanism driving the particular phase transition of the lattice gauge model under study.
@@
    LGW transitions, in which only gauge-invariant correlations of the matter field are critical. They admit a conventional effective LGW field-theoretical description in terms of a gauge-invariant order-parameter field, which orders at the transition, driving the spontaneous breaking of the global symmetry. Gauge fields do not display critical correlations. The gauge symmetry only prevents nongauge-invariant observables from becoming critical.
@@
    ${}^{\,\times}$ transitions, in which only matter fields develop critical correlations, similarly to LGW transitions. They also admit a LGW description, but in this case the order-parameter field is not gauge invariant. [^3] It is important to note that the gauge symmetry can never be broken [231, 232, 43, 8], despite the common use of referring to the Higgs mechanism as a spontaneous gauge symmetry breaking. The statement that in some cases a nongauge invariant order parameter is needed does not refer to the description of the thermodynamic phases, but to the definition of the effective model which encodes the universal critical properties of the continuous phase transition. In LGW× transitions the critical behavior of gauge-invariant quantities is the same as that of composite operators defined in terms of a nongauge invariant field, which is the fundamental field in the effective LGW description. As it is not gauge invariant, the order-parameter field can be identified in the gauge model only if an appropriate gauge fixing condition is added. Therefore, to distinguish them from standard LGW transitions, we add the superscript $\times$. [^4] Note that transitions with similar properties have been sometimes called LGW transitions, see, e.g., Ref. [21]. In particular, systems with discrete gauge groups undergo LGW× transitions for sufficiently weak gauge couplings.
@@
    GFT transitions, in which the critical behavior is due to a nontrivial interplay of scalar and gauge critical excitations. Therefore, their description requires an effective gauge field theory (GFT), with dynamical gauge and matter fields. GFT transitions are observed in the presence of continuous Abelian and non-Abelian gauge symmetries. They also occur in models that are locally invariant under discrete groups when the gauge symmetry enlarges at criticality, so that the critical modes are invariant under a continuous group.
@@
    Topological transitions, driven by topological modes, in which one cannot identify a local order-parameter field. A typical example is provided by the topological transition of the 3D lattice ${\mathbb{Z}}_{2}$-gauge model [120].